Differentiate the following w.r.t x by first principle
log(cosx)

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

(d/dx)log(cosx)=(d/dcosx)log(cosx)*(d/dx)cosx, first find dlog(cosx)/d(cosx)=f(x),let y=cos(x)

f(x)=limΔy→0(log(y+Δy)-log(y))/Δy=limΔy→0(log(1+ Δy/y)(1/Δy))=log(e(1/y))=1/y=1/cosx,d(cosx)/dx=

limΔx→0(cos(x+Δx)-cosx)/Δx =limΔx→0((cosx)cos(Δx)-(sinx)sin(Δx)-(cosx))/Δx =-sinx

therefore, answer is (1/cosx)(-sinx)=-tanx

rajula vineet reddy
12 years ago

use the chain rule i.e dy/dx=dy/dt * dt/dx .......dlog(cosx)/dx=dlog(cosx)/d(cosx) * d(cosx)/dx=1/(cosx) * -(sinx)=-(tanx)............similarly u can apply first principle to them using this rule................plz approve my answer if u like it